iimaniiii:

A text font fits 12 characters per inch. Using the same font, how many characters can be expected per yard of text? There are 12 inches in a foot, and 3 feet in a yard.

Extrinix:

Think of it like this, a yard is how many inches, well a yard is 36 inches. \(12in = 1ft\) \(3ft = 1yd\) Which would make it this: \(12in \times 3 = 1yd\) \(36 = 1yd\) If 12 characters fit in an inch, then you would multiply the amount of characters by how many inches, 36. \(amt~of~char~per~inch \times amt~of~inch~in~yd = amt~of~chars~per~yd\) \(12 \times 36 = ~?yd\)